the rear wheels of a car are turning at an angular speed of 60 rad/s. the brakes are applided for 5 s,causing a uniform angular retardation of 8 rads/s^2. the number of revolutionbs turned by the rear wheels during the braking period is about
 
 

Grade:9

1 Answers

Saurav
4 years ago
the basic fact to solve this question is 2pie radians make one revolution
angular distance covered by wheel be s
angular accleration be a= – 8 with unit
time be t =5s
initial angular velocity be u =60 again with unit
s=ut-(1/2)at*t
s=300-100=200radians
number of revoltions = 200/2pie
that is 100/pie
 
 
 
aprove if it helps you as it motivates me to answer more